# Trigonometry of partially truncated hyperbolic triangles and tetrahedra

### v2

@@
Trigonometry of partially truncated
hyperbolic triangles and tetrahedra
Jason DeBlois Department of Mathematics, University of Pittsburgh jdeblois@pitt.edu
For $n=4$, $5$ or $6$, an $n$-gon can be regarded as a partially truncated triangle, resulting from removing open affine neighborhoods of certain vertices of a triangle. This has particular significance in hyperbolic geometry when the edges resulting from truncation are at right angles to the others. Multiple versions of the hyperbolic laws of sines and cosines are proved for such polygons in Chapter VI of Fenchel’s Elementary geometry in hyperbolic space [7].
The first main results of this note, proved in Section 2, establish additional such laws which are not already in [7] nor other standard references that we know (eg. by Beardon [1] or Ratcliffe [10]). Proposition 2.1 pertains to hyperbolic quadrilaterals, and 2.2 to pentagons, in which the non-truncated vertices are ideal (see Section 1 below). In these results, the length of a horospherical cross-section at an ideal vertex plays the role that would be played by the dihedral angle at an ordinary vertex, and the distance to the cross-section plays the role of edge length.
In Sections 3 and 4 we turn to the analogous situation in dimension three. The results of Section 3 concern truncated tetrahedra in $\mathbb{H}^{3}$, which are homeomorphic to affine $3$-simplices with open neighborhoods of their vertices removed. Proposition 3.5 bounds the transversal length of such a tetrahedron, meaning the minimum distance between a specified pair of opposite internal edges, in terms of the tetrahedron’s set of internal edge lengths. Here an edge is internal if it is not contained in a face resulting from truncation. Section 4 addresses transversal lengths of partially truncated tetrahedra whose non-truncated vertices are ideal, successively treating the cases of one, two, three, and finally, in Proposition 4.8, four ideal vertices.
In all settings we follow the approach of [10, Ch. 3] and use the hyperboloid model for $\mathbb{H}^{n}$, where it is taken as a subset of $\mathbb{R}^{n+1}$ equipped with the Lorentzian inner product, a certain non positive-definite bilinear form. Vectors of the ambient $\mathbb{R}^{n+1}$ carry information about different objects of $\mathbb{H}^{n}$, depending on the sign of their self-pairing, and the Lorentzian inner product of two such vectors carries information about the hyperbolic distance between the objects the vectors encode. This property is exploited in [10] and also in previous work, eg. by Epstein–Penner [6]. We describe it in Section 1 and subsequently leverage it to encode partially truncated triangles using just three vectors, and truncated tetrahedra using just four, using the pairings of these vectors to prove our trigonometric formulas.
@@
The results of Section 3 here are applied in [5] as one tool contributing to that paper’s lower bounds on volumes of hyperbolic $3$-manifolds with totally geodesic boundary. Similarly, we expect the results of Sections 2 and 4 to be used in forthcoming work on non-compact such manifolds. We are not aware of other references in the literature for the main results of Sections 2, 3, and 4. Section 2 of [8] also considers the geometry of partially truncated tetrahedra, but tracks them using dihedral angles and is focused on different questions; eg. existence and moduli.
@@ 1. Background: the meaning of vectors in the hyperboloid model
It follows from Fact 1.1 that this formula is well-defined. It is the distance function determined by the Riemannian metric on $\mathbb{H}^{n}$ given, at each $\mathbf{x}\in\mathbb{H}^{n}$, by restricting the Lorentzian inner product to $T_{\mathbf{x}}\mathbb{H}^{n}=\mathbf{x}^{\perp}\doteq\{\mathbf{v}\,|\,\mathbf{v}\circ\mathbf{x}=0\}$. (This restriction is positive-definite since $\mathbf{x}$ is time-like, see [10, Theorem 3.1.5].) The isometry group of $\mathbb{H}^{n}$ is the group $O^{+}(1,n)$ of matrices preserving the Lorentzian inner product and the sign of time-like vectors, see [10, §3.1], acting on $\mathbb{H}^{n}$ by restriction.
@@ 1.1. The meaning of light-like vectors
Recall that $\mathbf{x}\in\mathbb{R}^{n+1}$ is light-like if $\mathbf{x}\circ\mathbf{x}=0$. Any positive light-like vector $\mathbf{x}$ is approached by a sequence of positive time-like vectors (for instance we can take $t\,\mathbf{x}+(1-t)\,\mathbf{e}_{1}$ for $t$ approaching $1$ from below); hence its projective class $[\mathbf{x}]$ in $\mathbb{R}P^{n}$ is approached by a sequence in the projectivization of $\mathbb{H}^{n}$. Conversely, the projectivization of the light cone is the frontier of the projectivization of $\mathbb{H}^{n}$ in $\mathbb{R}P^{n}$. In this sense we regard projectivized members of the light cone as ideal points of $\mathbb{H}^{n}$.
@@
The horosphere determined by a positive light-like vector $\mathbf{x}\in\mathbb{R}^{n+1}$ is $S=\{\mathbf{v}\in\mathbb{H}^{n}\,|\,\mathbf{v}\circ\mathbf{x}=-1\}$. The horoball bounded by $S$ is the set $B=\{\mathbf{v}\in\mathbb{H}^{n}\,|\,\mathbf{v}\circ\mathbf{x}\geq-1\}$. We say that the projective class $[\mathbf{x}]$ of $\mathbf{x}$ is the ideal point of $S$ or of $B$.
@@
$d_{S}(\mathbf{u}_{0},\mathbf{u}_{1})=\sqrt{-2(1+\mathbf{u}_{0}\circ\mathbf{u}_{1})}$
@@
For any $t\in\mathbb{R}$, $\gamma^{\prime}(t)=-\gamma(t)+e^{-d}e^{t}\mathbf{x}$ is a linear combination of $\gamma(t)$ and $\mathbf{x}$. Therefore by the discussion below Definition 1.2 it is normal to the horosphere through $\gamma(t)$ with ideal point $[\mathbf{x}]$.
@@
Therefore $e^{d_{H}(\mathbf{u},\mathbf{v})}$ is either $-\mathbf{v}\circ\mathbf{x}$ or its reciprocal, whichever is at least $1$ since $d_{H}(\mathbf{u},\mathbf{v})$ is non-negative. If we take $d$ to be the signed distance, with non-negative sign if $\mathbf{v}$ is outside the horoball $B$, then by the definition of $B$ we have $e^{d}=-\mathbf{v}\circ\mathbf{x}$ in all cases.
@@
For the parametrized curve $\gamma$ defined in the statement, direct computation reveals that $\gamma(t)\circ\gamma(t)=-1$ for all $t$, so $\gamma$ maps into $\mathbb{H}^{n}$, and that $\gamma^{\prime\prime}(t)=\gamma(t)$. Therefore $\gamma$ is a hyperbolic geodesic, by [10, Theorem 3.2.4]. More direct computation shows that $\gamma(0)=\mathbf{v}$ and $\gamma(d)$ is the nearest point $\mathbf{u}$ to $\mathbf{v}$ on $S$ described in (2). ∎
@@
A vector $\mathbf{u}\in\mathbb{R}^{n+1}$ lies in $S_{1}$ if and only if $\mathbf{u}\circ\mathbf{u}=-1$, $\mathbf{u}$ is positive, and $\mathbf{u}\circ\mathbf{x}_{1}=-1$. By the theory of Lagrange multipliers, the restriction of $f(\mathbf{u})=\mathbf{u}\circ\mathbf{x}_{0}$ to $B_{1}$ may attain a local extremum at $\mathbf{u}\in S_{1}$ only if the gradient of $f$ at $\mathbf{u}$ is a linear combination of the constraint gradients $\nabla g_{1}(\mathbf{u})$ and $\nabla g_{2}(\mathbf{u})$, where $g_{1}(\mathbf{u})=\mathbf{u}\circ\mathbf{x}_{1}$ and $g_{2}(\mathbf{u})=\mathbf{u}\circ\mathbf{u}$. Direct computation yields $\nabla f(\mathbf{u})=\bar{\mathbf{x}}_{0}$, $\nabla g_{1}(\mathbf{u})=\bar{\mathbf{x}}_{1}$, and $\nabla g_{2}(\mathbf{u})=2\bar{\mathbf{u}}$, where $\bar{\mathbf{x}}_{0}$ is obtained from $\mathbf{x}_{0}$ by multiplying the first entry by $-1$ and similarly for the others. It thus follows that at such a local extremum $\mathbf{u}$, $\mathbf{x}_{0}$ is a linear combination of $\mathbf{x}_{1}$ and $\mathbf{u}$ so, since $\mathbf{x}_{0}$ is not a multiple of $\mathbf{x}_{1}$, $\mathbf{u}$ is a linear combination of the $\mathbf{x}_{i}$.
@@ 1.2. The meaning of space-like vectors
As defined in [10, §3.2], a hyperplane of $\mathbb{H}^{n}$ is its intersection with a time-like, codimension-one vector subspace of $\mathbb{R}^{n+1}$. Corollary 4 of [10, §3.2] implies that the group of hyperbolic isometries acts transitively on the set of hyperplanes. Thus each hyperplane is the polar hyperplane to a space-like vector, since for instance $(\mathbb{R}^{n}\times\{0\})\cap\mathbb{H}^{n}$ is the polar hyperplane to $\mathbf{e}_{n+1}=(0,\ldots,0,1)$.
Every hyperplane $P=V\cap\mathbb{H}^{n}$ is a totally geodesic copy of $\mathbb{H}^{n-1}$ in $\mathbb{H}^{n}$, being, for any $\mathbf{x}\in P$, the image of the restriction of the exponential map based at $\mathbf{x}$ to $T_{\mathbf{x}}P=V\cap\mathbf{x}^{\perp}$. Conversely, the exponential map’s explicit description shows that any $(n-1)$-dimensional totally geodesic subspace $P$ of $\mathbb{H}^{n}$ is contained in $V=\mathrm{span}\{\mathbf{x},T_{\mathbf{x}}P\}$ for any $\mathbf{x}\in P$, and hence is a hyperplane.
We define a half-space to be the closure of one component of $\mathbb{H}^{n}-P$, for a hyperplane $P$. We call $P$ the boundary of $H$ and $H-P$ the interior. From eg. the model case above we see that each hyperplane bounds exactly two distinct half-spaces, which have disjoint interiors.
@@
Above, given a hyperplane $P$ and any $\mathbf{x}\in P$, a normal vector to $P$ at $\mathbf{x}$—and to a half-space $H$ bounded by $P$—is an element of $T_{\mathbf{x}}\mathbb{H}^{n}$ orthogonal to the codimension-one subspace $T_{\mathbf{x}}P$. A unit normal vector $\mathbf{y}$ to $P$ determines a geodesic $\gamma_{\mathbf{y}}(t)=\cosh t\,\mathbf{x}+\sinh t\,\mathbf{y}$ that intersects $P$ transversely, and we say $\mathbf{y}$ is outward to $H$ if $\gamma(t)\in H$ for all $t<0$.
@@
For a hyperplane $P$ and any $\mathbf{x}\in P$, since the orthogonal subspace to $T_{\mathbf{x}}P$ in $T_{\mathbf{x}}\mathbb{H}^{n}$ is one-dimensional there are exactly two unit normals to $P$. If $\mathbf{y}$ is one of these, the other is $-\mathbf{y}$, and exactly one of them is outward to a given half-space $H$ bounded by $P$. Take this to be $\mathbf{y}$. Any $\mathbf{x}^{\prime}\in P$ is of the form $\gamma_{\mathbf{z}}(1)$ for some $\mathbf{z}\in T_{\mathbf{x}}P$, with $\gamma_{\mathbf{z}}$ as in (1)—ie. $\mathbf{x}^{\prime}$ is the exponential image of $\mathbf{z}$—and hence $\mathbf{y}\circ\mathbf{x}^{\prime}$ also equals 00. Thus $P$ is the polar hyperplane of $\mathbf{y}$.
@@
A vector $\mathbf{v}\in\mathbb{R}^{n+1}$ lies in $P$ if and only if $\mathbf{v}\circ\mathbf{v}=-1$, $\mathbf{v}$ is positive, and $\mathbf{v}\circ\mathbf{y}=0$. By the theory of Lagrange multipliers, the restriction of $f(\mathbf{v})\doteq\mathbf{v}\circ\mathbf{x}$ to $P$ may attain a local extremum at $\mathbf{v}\in P$ only if the gradient of $f$ at $\mathbf{v}$ is a linear combination of the constraint gradients $\nabla g_{1}(\mathbf{v})$ and $\nabla g_{2}(\mathbf{v})$, where $g_{1}(\mathbf{v})=\mathbf{v}\circ\mathbf{y}$ and $g_{2}(\mathbf{v})=\mathbf{v}\circ\mathbf{v}$. Direct computation yields $\nabla f(\mathbf{v})=\bar{\mathbf{x}}$, $\nabla g_{1}(\mathbf{v})=\bar{\mathbf{y}}$, and $\nabla g_{2}(\mathbf{v})=2\bar{\mathbf{v}}$, where $\bar{\mathbf{x}}$ is obtained from $\mathbf{x}$ by multiplying the first entry by $-1$ and similarly for the others. It thus follows that $\mathbf{x}$ is a linear combination of $\mathbf{y}$ and $\mathbf{v}$ for such a point $\mathbf{v}$, so since $\mathbf{x}$ is not a multiple of $\mathbf{y}$ we can express $\mathbf{v}$ in terms of $\mathbf{x}$ and $\mathbf{y}$.
@@
where $\mathbf{x}_{0}\cdot\mathbf{y}_{0}$ is the ordinary dot product of $\mathbf{x}_{0}$ and $\mathbf{y}_{0}$. Since $\mathbf{x}\circ\mathbf{y}<0$ we must have $\mathbf{x}_{0}\cdot\mathbf{y}_{0}<0$; by the Cauchy Schwarz inequality, $-\mathbf{x}_{0}\cdot\mathbf{y}_{0}\leq\|\mathbf{x}_{0}\|\|\mathbf{y}_{0}\|$. Thus we have:
@@
In case (3) above we say that $P_{1}$ and $P_{2}$ are parallel. One can show in this case that there are sequences in $P_{1}$ and $P_{2}$ such that the infimum of distances from points of the first sequence to points of the second is 00. We now expand on case (2) above.
@@
where the sign of “$\pm$” above is negative if $\mathbf{v}_{1}$ belongs to the half-space $H_{2}$ bounded by $P_{2}$ with $\mathbf{y}_{2}$ as outward normal vector, and negative otherwise.
@@
By Lemma 1.6, $\mathbf{v}_{1}$ belongs to the half-space $H_{2}$ with $\mathbf{y}_{2}$ as outward normal if and only if $\mathbf{v}_{1}\circ\mathbf{y}_{2}<0$, hence if and only if the “$\pm$” above is negative.∎

### v3

@@
Abstract.
The first main results of this note establish forms of the hyperbolic laws of cosines and sines for certain classes of quadrilaterals and pentagons in the hyperbolic plane, having at least one ideal vertex and right angles at non-ideal vertices, in which the length of a horocyclic cross-section at an ideal vertex plays the role filled by the dihedral angle in the usual versions of these laws. The second set of main results concern transversal length, meaning the distance from a designated internal edge to its opposite, of partially truncated tetrahedra in three-dimensional hyperbolic space whose non-truncated vertices are ideal. Transversal lengths of such tetrahedra are proved to depend only on the entire collection of internal edge lengths (interpreted at ideal vertices in terms of horospherical cross-sections), and bounds on these lengths are established. The case of ideal tetrahedra (no truncated vertices) is also considered. All main results are established using the unifying perspective of the hyperboloid model and Lorentzian geometry. A thorough introduction to this perspective is provided, with references as appropriate.
This paper proves a family of trigonometric results about truncated and partially truncated hyperbolic triangles (in $\mathbb{H}^{2}$) and tetrahedra (in $\mathbb{H}^{3}$). Here in the two-dimensional setting we produce an $n$-gon, for $n=4$, 55 or $6$, by removing open neighborhoods of one, two, or all three vertices, respectively, of a triangle. Geometrically, we require each edge resulting from truncation to be at right angles to the others that it intersects. Similarly, a partially truncated tetrahedron is obtained by truncating some or all vertices of a tetrahedron; and geometrically, we require each triangular face resulting from truncation to intersect other faces at right angles.
The results of Section 4 give lower bounds, or in some cases formulas, for transversal lengths of partially truncated tetrahedra whose non-truncated vertices are ideal. A transversal length of such a tetrahedron is the minimum distance between a specified pair of opposite internal edges, meaning edges not produced by truncation. In order to capture the flavor of these results we restate the final one here, in slightly more accessible form. It pertains to (fully) ideal tetrahedra.
Proposition 4.8.
For an ideal tetrahedron $\Delta\subset\mathbb{H}^{3}$ and a choice of horoballs $B_{1},B_{2},B_{3},B_{4}$, one centered at each ideal vertex of $\Delta$, let $d_{ij}$ be the signed distance between $B_{i}$ and $B_{j}$ for each $i<j$ (with a negative sign if the horoballs overlap). If $\tilde{\lambda}_{12}$ is the geodesic joining the center of $B_{1}$ to that of $B_{2}$, and $\tilde{\lambda}_{34}$ is the geodesic joining the centers of $B_{3}$ to $B_{4}$, the length of the transversal of $\Delta$ joining $\tilde{\lambda}_{12}$ to $\tilde{\lambda}_{34}$ is given by
$\cosh T_{4}(x,y;a,b,c,d)=\frac{\sqrt{ad}+\sqrt{bc}}{\sqrt{xy}},$
where $x=e^{d_{12}}$, $y=e^{d_{34}}$, $a=e^{d_{13}}$, $b=e^{d_{14}}$, $c=e^{d_{23}}$, and $d=e^{d_{24}}$.
One sees from this that the transversal length depends only on the internal edge lengths, that it decreases with the length of either of the internal edges that it joins, and that it increases with the other four internal edge lengths. It is also invariant under swapping the index $B_{1}$ with $B_{2}$, and/or $B_{3}$ with $B_{4}$—these preserve the roles of $\tilde{\lambda}_{12}$ and $\tilde{\lambda}_{13}$ and act on the set $\{a,b,c,d\}$ of inputs by even involutions. Finally, it is invariant under changing the choice of horoball centered at any ideal vertex of $\Delta$, which rescales the distances along edges incident to that vertex, cf. Remarks 1. Thus although computing the transversal length requires choosing a set of horoballs, the length computed does not depend on the particular set chosen.
The prior results of Section 4, which address transversal lengths of partially truncated tetrahedra with one, two, or three ideal vertices, as well as those of Section 3 concerning fully truncated tetrahedra in $\mathbb{H}^{3}$, have the same set of qualitative properties. In these results, the length of an edge joining two faces resulting from truncation, or the distance from a truncation plane to a horoball, replace the signed distance between horoballs as appropriate. In some cases, like that of fully truncated tetrahedra, we give lower bounds rather than explicit formulas for the transversal length, due to the complexity of the formulas involved. See eg. Proposition 3.5.
@@ Abstract.
The results of Section 3 here are applied in [5] as one tool contributing to that paper’s lower bounds on volumes of hyperbolic $3$-manifolds with totally geodesic boundary, and we expect those of Sections 2 and 4 to be similarly useful in forthcoming work.
To prove these results we use the hyperboloid model for $\mathbb{H}^{n}$, where it is taken as a subset of $\mathbb{R}^{n+1}$ equipped with the Lorentzian inner product, a certain non positive-definite bilinear form. Vectors of the ambient $\mathbb{R}^{n+1}$ carry information about different objects of $\mathbb{H}^{n}$, depending on the sign of their self-pairing, and the Lorentzian inner product of two such vectors carries information about the hyperbolic distance between the objects the vectors encode. This property was exploited by Ratcliffe in [11, Ch. 3] to prove trigonometric formulas, and also in previous work, eg. by Epstein–Penner [6]. We describe it in Section 1 and subsequently leverage it to encode partially truncated triangles using just three vectors, and truncated tetrahedra using just four, using the pairings of these vectors to prove our trigonometric formulas.
In Section 2 we prove a set of hyperbolic trigonometric laws for partially truncated triangles in $\mathbb{H}^{2}$ that we do not know of in the standard references for such results, eg. written by Fenchel [7] (Chapter VI there has a vast collection), nor by Beardon [1] or Ratcliffe [11]. In the results that we prove, the length of a horospherical cross-section at an ideal vertex plays the role that would be played by the dihedral angle at an ordinary vertex, and the distance to the cross-section plays the role of edge length. Proposition 2.1 pertains to hyperbolic quadrilaterals, and 2.2 to pentagons, in which the non-truncated vertices are ideal (see Section 1 below). There is even a set of laws for ideal triangles, reproduced below, for which we do not know a full reference.
Proposition 2.3.
Let $\Delta$ be an ideal triangle in $\mathbb{H}^{2}$ and $B_{1},B_{2},B_{3}$ be horoballs, one centered at each ideal vertex of $\Delta$. For $i\in\{1,2,3\}$, let $\theta_{i}$ be the length of the horocyclic arc $S_{i}\cap\Delta$, where $S_{i}=\partial B_{i}$, and for each $i<j$ let $d_{ij}$ be the signed distance between $B_{i}$ and $B_{j}$ along the geodesic joining their centers (with a negative sign if the horoballs overlap). Then:
  • (Law of Sines) ${\frac{\theta_{1}}{e^{d_{23}}}=\frac{\theta_{2}}{e^{d_{13}}}=\frac{\theta_{3}}{e^{d_{12}}}}$
  • (First Law of Cosines) ${\theta_{1}=\sqrt{\frac{e^{d_{23}}}{e^{d_{12}}e^{d_{13}}}}}$
  • (Second Law of Cosines) ${e^{d_{23}}=\frac{1}{\theta_{2}\theta_{3}}}$ [This was proved as Lemma 3.3 of [9].]
The names above reference the corresponding results for compact hyperbolic triangles, see eg. [11, Theorems 3.5.2, 3.5.3, 3.5.4]. We are not aware of other references in the literature for the main results of Sections 2, 3, and 4, save for the “Second Law of Cosines” above. Section 2 of [8] also considers the geometry of partially truncated tetrahedra, but tracks them using dihedral angles and is focused on different questions; eg. existence and moduli.
@@ 1. Background: the meaning of vectors in the hyperboloid model
It follows from Fact 1.1 that this formula is well-defined. It is the distance function determined by the Riemannian metric on $\mathbb{H}^{n}$ given, at each $\mathbf{x}\in\mathbb{H}^{n}$, by restricting the Lorentzian inner product to $T_{\mathbf{x}}\mathbb{H}^{n}=\mathbf{x}^{\perp}\doteq\{\mathbf{v}\,|\,\mathbf{v}\circ\mathbf{x}=0\}$. (This restriction is positive-definite since $\mathbf{x}$ is time-like, see [11, Theorem 3.1.5].) The isometry group of $\mathbb{H}^{n}$ is the group $O^{+}(1,n)$ of matrices preserving the Lorentzian inner product and the sign of time-like vectors, see [11, §3.1], acting on $\mathbb{H}^{n}$ by restriction.
@@ 1.1. The meaning of light-like vectors
Recall that $\mathbf{x}\in\mathbb{R}^{n+1}$ is light-like if $\mathbf{x}\circ\mathbf{x}=0$. Any positive light-like vector $\mathbf{x}$ is approached by a sequence of positive time-like vectors (for instance we can take $t\,\mathbf{x}+(1-t)\,\mathbf{e}_{1}$ for $t$ approaching $1$ from below); hence its projective class $\mathbf{x}]$ in $\mathbb{R}P^{n}$ is approached by a sequence in the projectivization of $\mathbb{H}^{n}$. Conversely, the projectivization of the light cone is the frontier of the projectivization of $\mathbb{H}^{n}$ in $\mathbb{R}P^{n}$. In this sense we regard projectivized members of the light cone as ideal points of $\mathbb{H}^{n}$.
@@
The horosphere determined by a positive light-like vector $\mathbf{x}\in\mathbb{R}^{n+1}$ is $S=\{\mathbf{v}\in\mathbb{H}^{n}\,|\,\mathbf{v}\circ\mathbf{x}=-1\}$. The horoball bounded by $S$ is the set $B=\{\mathbf{v}\in\mathbb{H}^{n}\,|\,\mathbf{v}\circ\mathbf{x}\geq-1\}$. We say that the projective class $\mathbf{x}]$ of $\mathbf{x}$ is the ideal point of $S$ or of $B$.
@@
(2) $d_{S}(\mathbf{u}_{0},\mathbf{u}_{1})=\sqrt{-2(1+\mathbf{u}_{0}\circ\mathbf{u}_{1})}$
@@
For any $t\in\mathbb{R}$, $\gamma^{\prime}(t)=-\gamma(t)+e^{-d}e^{t}\mathbf{x}$ is a linear combination of $\gamma(t)$ and $\mathbf{x}$. Therefore by the discussion below Definition 1.2 it is normal to the horosphere through $\gamma(t)$ with ideal point $\mathbf{x}]$.
@@
Therefore $e^{d_{H}(\mathbf{u},\mathbf{v})}$ is either $\mathbf{v}\circ\mathbf{x}$ or its reciprocal, whichever is at least $1$ since $d_{H}(\mathbf{u},\mathbf{v})$ is non-negative. If we take $d$ to be the signed distance, with non-negative sign if $\mathbf{v}$ is outside the horoball $B$, then by the definition of $B$ we have $e^{d}=-\mathbf{v}\circ\mathbf{x}$ in all cases.
@@
For the parametrized curve $\gamma$ defined in the statement, direct computation reveals that $\gamma(t)\circ\gamma(t)=-1$ for all $t$, so $\gamma$ maps into $\mathbb{H}^{n}$, and that $\gamma^{\prime\prime}(t)=\gamma(t)$. Therefore $\gamma$ is a hyperbolic geodesic, by [11, Theorem 3.2.4]. More direct computation shows that $\gamma(0)=\mathbf{v}$ and $\gamma(d)$ is the nearest point $\mathbf{u}$ to $\mathbf{v}$ on $S$ described in (3). ∎
@@
A vector $\mathbf{u}\in\mathbb{R}^{n+1}$ lies in $S_{1}$ if and only if $\mathbf{u}\circ\mathbf{u}=-1$, $\mathbf{u}$ is positive, and $\mathbf{u}\circ\mathbf{x}_{1}=-1$. By the theory of Lagrange multipliers, the restriction of $f(\mathbf{u})=\mathbf{u}\circ\mathbf{x}_{0}$ to $B_{1}$ may attain a local extremum at $\mathbf{u}\in S_{1}$ only if the gradient of $f$ at $\mathbf{u}$ is a linear combination of the constraint gradients $\nabla g_{1}(\mathbf{u})$ and $\nabla g_{2}(\mathbf{u})$, where $g_{1}(\mathbf{u})=\mathbf{u}\circ\mathbf{x}_{1}$ and $g_{2}(\mathbf{u})=\mathbf{u}\circ\mathbf{u}$. Direct computation yields $\nabla f(\mathbf{u})=\bar{\mathbf{x}}_{0}$, $\nabla g_{1}(\mathbf{u})=\bar{\mathbf{x}}_{1}$, and $\nabla g_{2}(\mathbf{u})=2\bar{\mathbf{u}}$, where $\bar{\mathbf{x}}_{0}$ is obtained from $\mathbf{x}_{0}$ by multiplying the first entry by 1-1 and similarly for the others. It thus follows that at such a local extremum $\mathbf{u}$, $\mathbf{x}_{0}$ is a linear combination of $\mathbf{x}_{1}$ and $\mathbf{u}$ so, since $\mathbf{x}_{0}$ is not a multiple of $\mathbf{x}_{1}$, $\mathbf{u}$ is a linear combination of the $\mathbf{x}_{i}$.
@@ 1.2. The meaning of space-like vectors
As defined in [11, §3.2], a hyperplane of $\mathbb{H}^{n}$ is its intersection with a time-like, codimension-one vector subspace of $\mathbb{R}^{n+1}$. Corollary 4 of [11, §3.2] implies that the group of hyperbolic isometries acts transitively on the set of hyperplanes. Thus each hyperplane is the polar hyperplane to a space-like vector, since for instance $(\mathbb{R}^{n}\times\{0\})\cap\mathbb{H}^{n}$ is the polar hyperplane to $\mathbf{e}_{n+1}=(0,\ldots,0,1)$.
Every hyperplane $P=V\cap\mathbb{H}^{n}$ is a totally geodesic copy of $\mathbb{H}^{n-1}$ in $\mathbb{H}^{n}$, being, for any $\mathbf{x}\in P$, the image of the restriction of the exponential map based at $\mathbf{x}$ to $T_{\mathbf{x}}P=V\cap\mathbf{x}^{\perp}$. Conversely, the exponential map’s explicit description shows that any (n1)(n-1)-dimensional totally geodesic subspace $P$ of $\mathbb{H}^{n}$ is contained in $V=\mathrm{span}\{\mathbf{x},T_{\mathbf{x}}P\}$ for any $\mathbf{x}\in P$, and hence is a hyperplane.
We define a half-space to be the closure of one component of $\mathbb{H}^{n}-P$, for a hyperplane $P$. We call $P$ the boundary of $H$ and HPH-P the interior. From eg. the model case above we see that each hyperplane bounds exactly two distinct half-spaces, which have disjoint interiors.
@@
Above, given a hyperplane $P$ and any $\mathbf{x}\in P$, a normal vector to $P$ at $\mathbf{x}$—and to a half-space $H$ bounded by PP—is an element of $T_{\mathbf{x}}\mathbb{H}^{n}$ orthogonal to the codimension-one subspace $T_{\mathbf{x}}P$. A unit normal vector $\mathbf{y}$ to $P$ determines a geodesic $\gamma_{\mathbf{y}}(t)=\cosh t\,\mathbf{x}+\sinh t\,\mathbf{y}$ that intersects $P$ transversely, and we say $\mathbf{y}$ is outward to $H$ if $\gamma(t)\in H$ for all $t<0$.
@@
For a hyperplane $P$ and any $\mathbf{x}\in P$, since the orthogonal subspace to $T_{\mathbf{x}}P$ in $T_{\mathbf{x}}\mathbb{H}^{n}$ is one-dimensional there are exactly two unit normals to $P$. If $\mathbf{y}$ is one of these, the other is $\mathbf{y}$, and exactly one of them is outward to a given half-space $H$ bounded by $P$. Take this to be $\mathbf{y}$. Any $\mathbf{x}^{\prime}\in P$ is of the form $\gamma_{\mathbf{z}}(1)$ for some $\mathbf{z}\in T_{\mathbf{x}}P$, with $\gamma_{\mathbf{z}}$ as in (1)—ie. $\mathbf{x}^{\prime}$ is the exponential image of $\mathbf{z}$—and hence $\mathbf{y}\circ\mathbf{x}^{\prime}$ also equals 0. Thus $P$ is the polar hyperplane of $\mathbf{y}$.
@@
A vector $\mathbf{v}\in\mathbb{R}^{n+1}$ lies in $P$ if and only if $\mathbf{v}\circ\mathbf{v}=-1$, $\mathbf{v}$ is positive, and $\mathbf{v}\circ\mathbf{y}=0$. By the theory of Lagrange multipliers, the restriction of $f(\mathbf{v})\doteq\mathbf{v}\circ\mathbf{x}$ to $P$ may attain a local extremum at $\mathbf{v}\in P$ only if the gradient of $f$ at $\mathbf{v}$ is a linear combination of the constraint gradients $\nabla g_{1}(\mathbf{v})$ and $\nabla g_{2}(\mathbf{v})$, where $g_{1}(\mathbf{v})=\mathbf{v}\circ\mathbf{y}$ and $g_{2}(\mathbf{v})=\mathbf{v}\circ\mathbf{v}$. Direct computation yields $\nabla f(\mathbf{v})=\bar{\mathbf{x}}$, $\nabla g_{1}(\mathbf{v})=\bar{\mathbf{y}}$, and $\nabla g_{2}(\mathbf{v})=2\bar{\mathbf{v}}$, where $\bar{\mathbf{x}}$ is obtained from $\mathbf{x}$ by multiplying the first entry by 1-1 and similarly for the others. It thus follows that $\mathbf{x}$ is a linear combination of $\mathbf{y}$ and $\mathbf{v}$ for such a point $\mathbf{v}$, so since $\mathbf{x}$ is not a multiple of $\mathbf{y}$ we can express $\mathbf{v}$ in terms of $\mathbf{x}$ and $\mathbf{y}$.
@@
where $\mathbf{x}_{0}\cdot\mathbf{y}_{0}$ is the ordinary dot product of $\mathbf{x}_{0}$ and $\mathbf{y}_{0}$. Since $\mathbf{x}\circ\mathbf{y}<0$ we must have $\mathbf{x}_{0}\cdot\mathbf{y}_{0}<0$; by the Cauchy Schwarz inequality, $\mathbf{x}_{0}\cdot\mathbf{y}_{0}\leq\|\mathbf{x}_{0}\|\|\mathbf{y}_{0}\|$. Thus we have:
@@
In case (3) above we say that $P_{1}$ and $P_{2}$ are parallel. One can show in this case that there are sequences in $P_{1}$ and $P_{2}$ such that the infimum of distances from points of the first sequence to points of the second is 0. We now expand on case (2) above.
@@
where the sign of $\pm$” above is negative if $\mathbf{v}_{1}$ belongs to the half-space $H_{2}$ bounded by $P_{2}$ with $\mathbf{y}_{2}$ as outward normal vector, and positive otherwise.
@@
By Lemma 1.6, $\mathbf{v}_{1}$ belongs to the half-space $H_{2}$ with $\mathbf{y}_{2}$ as outward normal if and only if $\mathbf{v}_{1}\circ\mathbf{y}_{2}<0$, hence if and only if the $\pm$” above is negative.∎